find the minor axis of the ellipse using the givin equation [(x+4)^2 over (6^2)] + [(y+8)^2 over (7^2)]=1
Answer by lwsshak3(11628)

find the minor axis of the ellipse using the givin equation [(x+4)^2 over (6^2)] + [(y+8)^2 over (7^2)]=1
**
(x+4)^2/(6^2) + (y+8)^2/(7^2)=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1 , a>b, with (h,k) being the (x,y) coordinates of the center
For given equation:
center: (-4,-8)
a=7
b=6
length of minor axis=2b=12
end points of minor axis: (-4±b,-8)=(-4±6,-8) or (-10,-8) and (2,-8)